# DreamStory: Open-Domain Story Visualization by LLM-Guided Multi-Subject Consistent Diffusion

### v2

@@ 4.5.1 Objective Comparison
All the objective results are presented in Tab. III. It is evident from Tab. III that adding MMSA and MMCA improved subject consistency, as indicated by an increase in DreamSim similarity (DS) and D&C-DS. However, a minor decline was observed in the similarity between the scene and its text. This is attributed to our generation process’s emphasis on subject consistency, which marginally affects the scene’s content. Nonetheless, this impact is negligible, i.e., the difference in CLIP-T similarity is less than 0.007. These results confirm that our method maintains subject consistency while preserving the scene’s semantics. Moreover, Tab. III reveals that incorporating both MMSA and MMCA modules led to our DreamStory achieving optimal performance regarding aesthetic scores and subject consistency. This conclusively validates the effectiveness of our approach.
@@
4.5.2 Subjective Comparison
We also conducted an ablation study about rewriting and showed the result in Fig. 15. As illustrated in Fig. 15, without rewriting, the diffusion model fails to understand the names, such as Kondo. This leads to generating incorrect types of subjects, such as generating gorillas as humans, as shown in the second row of Fig. 15.

### v3

@@ 4.5.1 Objective Comparison
All the objective results are presented in Tab. III. It is evident from Tab. III that adding MMSA and MMCA improved subject consistency, as indicated by an increase in DreamSim similarity (DS) and D&C-DS. However, a minor decline was observed in the similarity between the scene and its text. Quantitative analysis reveals that the CLIP-T score degradation primarily stems from background regions. The average CLIP-T score of background regions (masked subjects with 0/255 values) showed markedly greater reduction compared to complete scenes. Specifically, in the 3-subject benchmark, the background CLIP-T score drops by -0.0100, while the full scene declines only by -0.0050. A similar trend holds for the 2-subject benchmark, where the background decreases by -0.0051, compared to -0.0039 for the full scene. This phenomenon is attributed to our generation process’s emphasis on subject consistency, which marginally affects the scene’s content. Nonetheless, this impact is negligible, i.e., the difference in CLIP-T similarity is less than 0.007. These results confirm that our method maintains subject consistency while preserving the scene’s semantics. Moreover, Tab. III reveals that incorporating both MMSA and MMCA modules led to our DreamStory achieving optimal performance regarding aesthetic scores and subject consistency. This conclusively validates the effectiveness of our approach.
4.5.2 Subjective Comparison
We conducted an ablation study to evaluate the necessity of (1) LLM rewriting and (2) accurate mask generation. As illustrated in Fig. 14, without rewriting, the diffusion model fails to understand the names, such as Kondo. This leads to generating incorrect types of subjects, such as generating gorillas as humans, as shown in the second row of Fig. 14. Similarly, without our accurate mask generation strategy, the initial mask fails to fully cover the boy when the generated boy exceeds the initial mask region. This leads to inconsistent color generation in the hair region, as highlighted with the red circle in Fig. 15. These results clearly validate the effectiveness of our proposed rewriting and mask generation approach.
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (U1911203, U2001211, U22B2060), Guangdong Basic and Applied Basic Research Foundation (2019B1515130001), Key-Area Research and Development Program of Guangdong Province (2020B0101100001, 2024B0101050005), Research Foundation of Science and Technology Plan Project of Guangzhou City (2023B01J0001, 2024B01W0004).